OpenStudy (anonymous):

Write 0.14 as a fraction in lowest terms. Enter your answer as a reduced fraction (like 5/3, not 10/6) or as a whole number (like 4 or -2)

OpenStudy (anonymous):

.xx = xx/100

OpenStudy (anonymous):

Tomo that doesn't help me

OpenStudy (anonymous):

Okay, anything that's 0.xxxxx is less than one, so we won't have a problem with numbers like 5/3 or the like. But let's start off by expressing the ".14" as a fraction.

jimthompson5910 (jim_thompson5910):

Similar Example: 0.36 = 36/100 reduce to get 36/100 = 18/50 = 9/25 So 0.36 = 9/25 This is NOT your answer as its just an example

OpenStudy (anonymous):

so how we make .14 as a fraction

OpenStudy (anonymous):

is it 14/100

OpenStudy (anonymous):

Look at Jim's example just above me. He uses .36, you use .14. Think.

OpenStudy (anonymous):

Yes, it's 14/100 !! Now we have to reduce that to it's lowest terms, again using Jim's example as a guideline.

OpenStudy (anonymous):

Are you still there?